OpenStudy (anonymous):

Differentiate \[\ln \sin^2 x\]

OpenStudy (anonymous):

take sin^2x = t. dt/dx = 2sinx cosx = sin2x now for y=lnt dy/dx = 1/t dt/dx = 1/sin^2x . 2sinx cosx = cosx/sinx =cotx

OpenStudy (anonymous):

sorry its 2cotx.

OpenStudy (jamesj):

Like your last problem, it is work you using some properties of logarithms to simply the problem. As \[ \ln y^2 = 2 \ln y, \] it follows that \[ \ln \sin^2 x = 2 \ln \sin x \] Now this function is a little easier to differentiate.

OpenStudy (anonymous):

thanks